find the least common multiple of 10, 15, and 25
Answer by stanbon(75887)

find the least common multiple of 10, 15, and 25
---
Factor:
10 = 2*5
15 = 3^5
25 = 5^2
--------------
The lcm must have each different factor in its highest power:
---------
lcm = 2*3*5^2 = 150
=======================
Cheers,
Stan H.